OpenStudy (anonymous):

solve z + 10 divided by 9 = 2

OpenStudy (saifoo.khan):

\[\frac{z+10}{9}=2\]

OpenStudy (saifoo.khan):

Like that?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

z=8

OpenStudy (karatechopper):

z=8

OpenStudy (anonymous):

z+10 =18 ===> z=8

OpenStudy (anonymous):

z+10 = 9(2) z=18-10 z=8